Loli:

Which expression is equivalent to (3r/s)^3(s^2/9) A~ r^3/s B~ 3r^3/s C~ 3r^3s D~ r^3s

Mehek:

Does it look like this \(\bf (\dfrac{3r}{s})^3(\dfrac{s^2}{9})\)

Loli:

Yep

Mehek:

Ok so let's apply the ^3 \((3r)^3=27r^3\) and \(s^3\) \(\bf\dfrac{27r^3}{s^3}\)

Mehek:

\(\bf (\dfrac{27r^3}{s^3})(\dfrac{s^2}{9})\) You can cross out s^2 from both s's which leaves you with \(\bf \dfrac{27r^3}{s}~and~\dfrac{1}{9}\) and then you can divide \(\bf 27r^3~and~9~by~3\) Which basically leaves you with \(\bf \dfrac{3r^3}{s}*\dfrac{1}{1}=\dfrac{3r^3}{s}\) Option B)

Loli:

Thank you O: ♥

Mehek:

No problem :)